Pat travels 70 miles on her milk route, and Bob travels 75 miles on his route. Pattravels 5 miles per hour slower than Bob, and her route takes her one-half hour longer than Bob's. How fast is each one traveling?
Answer by stanbon(48558)

Pat travels 70 miles on her milk route, and Bob travels 75 miles on his route. Pat travels 5 miles per hour slower than Bob, and her route takes her one-half hour longer than Bob's. How fast is each one traveling?
--------------------
Pat DATA:
distance = 70 miles; rate = x-5 755mph ; time = d/r = 70/(x-5) hrs.
----------
Bob DATA:
distance = 75 miles ; rate= x mph ; time = d/r = 75/x hrs
-------
EQUATION:
Bob rate + 0.5 = Pat rate
75/x + 1/2 = 70/(x-5)
(151/2x) = 70/(x-5)
140x = 151x-755
-11x = -755
x=68.64 mph (Bob's rate)
x-5 = 63.64 mph (Pat's rate)
-----------
Cheers,
Stan H.